# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

### 29 Nov 2023 v1

@@ 3.2 Thermal activations for super-resolution
The multi-scale signal representation defined above is a natural match for super-resolution, and particularly well-suited if arbitrary scale factors shall be covered. Still, two challenges must be addressed. First, our formulation restricts the choice of architecture. Second, while it theoretically guarantees the downsampling operation ($t>t_{0}$), the upsampling operation ($0<t<t_{0}$) remains ill-posed. To narrow down the (infinite) solution space to a unique result a prior must either be defined in an unsupervised fashion or learned from data. We describe below how these issues can be addressed, so that Thera can be used to model local fields that represent each pixel in the low-resolution image, a strategy initially proposed by LIIF [11] and used by most recent ASSR methods [22, 8, 41, 10].
@@
Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, where $N$ is the number of parameters in a field of the desired size. Given a low-resolution image, the hypernetwork maps it to parameters for local fields, each spanning the area of one pixel of the low-resolution input. Even though the fields themselves model only a local part of the image, the hypernetwork informs them with rich contextual features collected over a large receptive field. The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [11, 22, 41, 8, 10], followed by one or more layers which produce feature maps of depth $N$. In this work, we use a series of ConvNeXt [25] blocks for this purpose. The local fields themselves are then supervised with values from high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and temporal coordinate $t$. The latter ensures that the signal has been filtered to the correct target resolution prior to querying discrete pixel values. Similar to LIIF, we use a local coordinate system for every field that is bounded in $[-0.5,0.5]\times[-0.5,0.5]$. The components $\textbf{W}_{1}$ are optimized directly rather than being an output of $\Psi$. Not only does this better fit the idea of having a consistent basis with which we represent each signal, but it also reduces the number of total parameters and, as we show later, leads to better results.
@@
Figure 4: Overview of Thera. A hypernetwork estimates the parameter of local fields. Each field uses thermal activations for multi-scale signal representation and anti-aliasing.
@@
The described scheme, Thera, is depicted in Fig. 4. It allows for arbitrary-scale super-resolution combining the power of proven feature extractors for SR and image restoration with the presented multiscale representation capabilities. As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors from data that lie within the resolution range observed during training. E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $\frac{1}{16}<t$. By training using multiple resolutions, we can also make $\kappa$ a trainable parameter to allow the network to adapt to different downsampling operators. We also manually set the bias of each channel of $\Phi$ to the RGB values of its associated low-resolution pixel.
@@
Priors at $\bm{t=0}$. To perform ASSR for out-of-domain scaling factors, we can use unsupervised priors at $t=0$. Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods. In our implementation, that prior takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well. We use an $\ell^{1}$ variant of TV:
@@
Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically using automatic differentiation tools, rather than using a neighborhood approximation as done most often in previous work [36].
Implementation and training. We use JAX [7] to implement Thera to maximize the efficiency of the hypernetwork operations. A PyTorch [34] implementation to fit a Thera field to a single image is also provided. Similar to prior work [11, 22], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to a patch of size $48^{2}$. As corresponding targets, $48^{2}$ random pixels are sampled from the target patch. We train with the Adam optimizer [21] with an initial learning rate of $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$ for $5\times 10^{6}$ iterations with a batch size of 16. Throughout the optimization, the learning rate is decayed to zero according to a cosine annealing schedule [26]. We use a mean absolute error (MAE) loss for training, to which the TV loss from Eq. 7 is added with a weight of $10^{-4}$. Standard data augmentations are used during training, including random flipping, rotations, and resizing of images. Analogous to previous work [40, 24, 41], we implement geometric self-ensembling, where the results of four rotated versions of the input are averaged at test time. We did not observe an improvement in performance if flips were included. Our code and pre-trained checkpoints will be made publicly available.
@@ 4 Results
EDSR- baseline [24] (1.22 M) MetaSR [17] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
@@
We evaluate the performance of Thera on a variety of datasets and with three popular feature extractors as backbones. We also conduct an ablation study that supports the design choices described in Sec. 3. Throughout this section, we will evaluate three variants of our method which differ solely in size of the hypernetwork and field:
  • Thera-S: We set the number of shared components to 32, and omit all but one layer of the hypernetwork, the layer that maps features to field parameters. This version adds only 8,256 parameters on top of the backbone.
  • Thera-M: This version uses 512 shared components and a moderately sized hypernetwork (5 ConvNeXt layers), approximately matching the parameter count of medium-sized competing methods (MetaSR [17], LIIF [11], LTE [22] and CUF [41]). This amounts to 0.48 M additional parameters.
  • Thera-L: We use 512 components and a larger hypernetwork (16 ConvNeXt layers), amounting to 1.41 M additional parameters. This version is comparable to CiaoSR [8] in terms of parameter count.
Datasets and metrics. Like previous methods, our models are trained with the DIV2K [1] training set, which is composed of 800 high-resolution RGB images of diverse scenes. We report evaluation metrics on the official DIV2K validation split (100 images) as well as on a set of standard benchmark datasets: Set5 [6], Set14 [47], BSDS100 [27] (all three containing natural images of low to medium resolution), Urban100 [18] (high-resolution urban scenes), and Manga109 [28] (high-resolution Japanese manga illustrations). Following prior work, we use peak-signal-to-noise ratio (PSNR, in decibels) as the evaluation metric, where PSNR is computed in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets. Not all numbers could be computed for competing methods due to the lack of availability of code or checkpoints, see Appendix E. Additional quantitative metrics are also provided in the supplementary material.
Backbones. Following related work, we train our method jointly with three standard backbone feature extractors for super-resolution and image restoration: (i) EDSR-baseline [24] (1.22 M parameters); (ii) RDN [48] (22.0 M parameters); and (iii) SwinIR [23] (12.0 M parameters).
@@ 4.1 Super-resolution performance
We begin with evaluating the three variants of our method (Thera-S, -M, and -L) on the hold-out DIV2K validation set, following the setup described above. Table 1 shows the PSNR values obtained by all investigated methods, for both in-distribution (×$\times 2$×$\times 4$) and out-of-distribution (×$\times 6$×$\times 30$) scaling factors. Thera achieves state-of-the-art performance. Thera-L outperforms all previous methods on 22 out of 24 cases, often by a considerable margin (e.g., $\times$3 using EDSR-baseline). Thera-M, which requires only a third as many additional parameters, still outperforms or – less frequently – is on par with competitors that have comparable parameter count in most cases. Notably, except for ×$\times 2$, Thera-M also matches the performance of more expensive, attention-based methods CiaoSR and CLIT, despite a more than 30×30$\times$30 × fewer parameters than CLIT.
Notably, our minimal version Thera-S – enabling arbitrary-scale SR with theoretical guarantees using only 8,192 additional parameters – regularly outperforms all comparison methods except for the most parameter-intensive ones when using the EDSR backbone. Using the more powerful SwinIR backbone, Thera-S outperforms all previous methods (including attention-based ones) on 6 of 8 scaling factors, despite adding almost 2000×2000$\times$2000 × fewer learnable parameters than CLIT (and 170×170$\times$170 × fewer than CiaoSR) on top of the feature extractor. Similarly, with RDN our smallest variant performs best in 7 of 8 scaling factors. This confirms that having formal modeling of multi-scale representation baked into our network is highly beneficial, especially for higher scaling factors.
@@
We further report the performance of our method on the five considered benchmark datasets in Tab. 2. For these experiments, we use the SwinIR backbone. Overall, due to the out-of-distribution nature of the data and the small size of some legacy datasets (e.g., 5 small images for Set5), the results are less consistent. This suggests that future work on ASSR would benefit from using larger-scale datasets in the future. Nonetheless, our methods perform best in 12 out of 15 cases. We once again observe that the performance of Thera-S is often comparable to that of methods that require orders of magnitude more parameters.
In our qualitative analysis – an excerpt of which is given in Fig. 5 – we observe that Thera achieves a more faithful and perceptually more convincing reconstruction, in particular with repeating structure. This once again highlights the advantages of using an accurate observation model in our field formulation.
@@
Input (low res.) LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) Ground truth
Figure 5: Qualitative examples from four datasets for $\times$6 super-resolution. We use the SwinIR [23] backbone for all methods. More examples are given in Appendix B.
@@ 4.2 Multi-scale consistency
We are interested in learning a super-resolution model that is consistent across scales. For instance, if we super-resolve an image with a scaling factor of $\times$4 then downsample it by a factor of $\times$2, we expect the result to be very similar to what we get by simply super-resolving the original image by a factor of $\times$2. We report in Tab. 3 the quality of this match for several scaling factor pairs for Thera – which uses a formal observation model – and for three comparison methods that learn such model from data. First, we notice that better super-resolution quality does not correlate with scale consistency – e.g., LTE achieves better consistency than CiaoSR for most scaling factor pairs. Second, while other methods generally exhibit good consistency, even for out of distribution (OOD) scales, our formal model is superior for all scaling factor combinations, often by a large margin, while requiring only a single parameter $\kappa$ to be learned (if any, since $\kappa$ can also be fixed as we show below).
@@
Table 3: Comparison of multi-scale consistency (in PSNR) for various combinations of scaling factors.
@@ 4.3 Ablation studies
In Tab. 4, we ablate individual components and design choices of our method to understand their contribution to the overall performance. These comparisons are done using Thera-M using the EDSR-baseline backbone.
Single scale training. We ran three experiments using a single scale ($\times$2, $\times$3, $\times$4) to test how this affects scale generalization. $\kappa$ was fixed for these experiments since it requires multiscale training to be optimized. We observe equal or even superior performance compared to our base case for the scale used for training (marked in yellow in Tab. 4), but we see a significant drop in generalization.
Trainable $\kappa$. We observe a small drop in performance when using the fixed, theoretically derived value for $\kappa$ (Eq. 5). This suggests that there are still effects that have not been accounted for in our model, but the effects are very minor.
Geometric self-ensemble. As observed by previous work [40, 24, 41], we see a notable boost in performance when using geometric self-ensemble (GSE). Nevertheless, if a user prioritizes speed over quality, this addition can be disabled without retraining the network at test-time.
Total variation. We observe negligible effects for in-domain scaling factors when no TV is used, but performance degrades significantly for out-of-distribution scales.
No thermal activations. We replace thermal activations (Eq. 4) for standard ReLU activations. What remains is only the hypernetwork controlling the parameters of the local fields. A loss in performance shows the value of our proposed thermal activation for multiscale representation.
Shared components. We observe a noticeable performance drop in all scales if we make the components $\mathbf{W}_{1}$ an output of the hypernetwork $\Psi$. Despite being even more flexible, it seems challenging to learn per-field components and other correlated parameters simultaneously.
@@
Table 4: Ablations have been run using Thera-M using the EDSR-baseline backbone, a representative case for all other settings.
@@ 4.4 Limitations and future work
Thermal activations as introduced in this paper, and by extension our proposed ASSR method, come with relatively strict architectural requirements that only allow for a single hidden layer in the neural field. While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder. While we have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work. This could result in higher parameter efficiency, and potentially better generalization. There may also be other formulations with similar characteristics which do not require the usage of a hypernetwork, which may allow for more streamlined training.
We noticed a relative drop in performance w.r.t. other methods for $\times$2 SR. We speculate that our weakness relative to higher scaling factors is due to this scale being close to the interface between continuous (field) and grid (pixel) representations, which Thera’s formalism does not model. Further exploration on this transition regime for such hybrid structures would be useful to better understand and use these hybrid representations. This could also benefit Instant-NGP [31] and related methods.
Finally, we have only explored TV as a method for regularizing $\Phi$ at $t=0$, but we expect other signal priors to be even more effective. Priors at $t=0$, possible only by using Thera, have the potential to regularize the continuous signal itself, and therefore improve SR performance for all scaling factors.

### 14 Mar 2024 v2

@@ 3.2 Thermal Activations for Super-Resolution
The multi-scale signal representation defined above is a natural match for super-resolution, and particularly well-suited if arbitrary scale factors shall be covered. Still, two challenges must be addressed. First, our formulation restricts the choice of architecture to MLPs with a single hidden layer. Second, while it theoretically guarantees the downsampling operation ($t>t_{0}$), the upsampling operation ($0<t<t_{0}$) remains ill-posed. To narrow down the (infinite) solution space to a unique result, a prior must either be defined in an unsupervised fashion or learned from data.
@@
Our solution to both challenges is to condition local fields with a hypernetwork $\Psi:\mathbb{R}^{W\times H\times 3}\to\mathbb{R}^{W\times H\times N}$, where $N$ is the number of parameters in a field of the desired size. The hypernetwork maps the low-resolution input image to parameters of local fields, each spanning the area of one pixel of the low-resolution input as originally proposed by LIIF [11] and adopted by most recent ASSR methods [22, 8, 43, 10]. Even though the fields themselves model only a local part of the image, the hypernetwork informs them with rich contextual features collected over a large receptive field. The hypernetwork $\Psi$ is composed of a standard SR backbone, as used in previous work [11, 22, 43, 8, 10], followed by one or more layers that produce feature maps of depth $N$. In this work, we use a series of ConvNeXt [26] blocks. The local fields themselves are then supervised with values from high-resolution target pixels at the appropriate spatial coordinates $\mathbf{x}$ and time index $t$. The latter ensures that the signal has been filtered to the correct target resolution prior to querying discrete pixel values. Similar to LIIF, the local coordinate system for every field is bounded by $[-0.5,0.5]\times[-0.5,0.5]$. In practice, we directly optimize a single global frequency bank $\textbf{W}_{1}$, rather than having the hypernetwork predict a separate $\textbf{W}_{1}$ for each low-resolution pixel. Not only does this better fit the idea of to represent the signal with a single, consistent basis; it also reduces the total parameter count and empirically leads to better results.
@@
Figure 4: Overview of Thera. A hypernetwork estimates the parameters of local fields. Each field uses thermal activations for multi-scale signal representation and anti-aliasing.
@@
The described scheme, which we call Thera, is depicted in Fig. 4. It allows for arbitrary-scale super-resolution, combining the proposed multi-scale signal representation with the expressivity of proven feature extraction backbones for SR and image restoration. As the entire network is trained end-to-end, the feature extractor can learn super-resolution priors for a whole range of resolutions covered by the training data. E.g., a network trained with scaling factors up to $\times$4 will encode priors that enable us to observe the field at $\frac{1}{16}<t$. By training on multiple resolutions, we can also make $\kappa$ a trainable parameter that allows the network to adapt to different downsampling operators. Finally, we set the bias terms for the three color channels of every local field $\Phi$ (i.e., $\textbf{b}_{2}$) to the RGB values of the associated low-resolution pixel. Thus, the hypernetwork only predicts field-wise phase shifts $\mathbf{b}_{1}$ and amplitudes $\mathbf{W}_{2}$.
@@
Prior at $\bm{t=0}$. To perform ASSR for out-of-domain scaling factors, we can use an unsupervised prior at $t=0$. Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods. In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well. We use an $\ell^{1}$ variant of TV:
@@
Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically using automatic differentiation tools, rather than falling back to a neighborhood approximation as in most previous work [38].
Implementation and training. We have implemented Thera in JAX [7] to maximize the efficiency of the hypernetwork operations. A PyTorch [36] implementation to fit a Thera field to a single image is also provided. Similar to prior work [11, 22], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$. As corresponding targets, $48^{2}$ random pixels are sampled from the target patch. We train with the Adam optimizer [21] with batch size $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The learning rate is decayed to zero according to a cosine annealing schedule [27]. As reconstruction loss, we minimize the mean absolute error (MAE), to which the TV loss from Eq. 7 is added with a weight of $10^{-4}$. Training data is augmented by standard data augmentations, namely random flipping, rotation, and resizing. As in previous work [42, 24, 43], we implement geometric self-ensembling, where the results for four rotated versions of the input are averaged at test time. We did not observe a performance gain when also including reflections.
@@ 4 Results
We evaluate Thera on a variety of datasets and with three popular feature extractors as backbones. We also conduct an ablation study that supports the design choices described in Sec. 3. Throughout this section, we will evaluate three variants of our method which differ solely in the size of the hypernetwork and the number of field parameters:
  • Thera-S: A tiny version with the number of globally shared components in $\mathbf{W}_{1}$ set to 32, and with only a single layer in hypernetwork, which maps features to field parameters. This version adds only 8,256 parameters on top of the backbone.
  • Thera-M: This version uses 512 shared components and a moderately sized hypernetwork with 5 ConvNeXt layers. Its parameter count of 0.48 M approximately matches that of medium-sized competitors (MetaSR [17], LIIF [11], LTE [22] and CUF [43]).
  • Thera-L: A large version with 512 components and a deeper hypernetwork of 16 ConvNeXt layers, for a total of 1.41 M additional parameters, comparable to CiaoSR [8].
Datasets and metrics. Following previous work, our models are trained with the DIV2K [1] training set, consisting of 800 high-resolution RGB images of diverse scenes. We report evaluation metrics on the official DIV2K validation split (100 images) as well as on a set of standard benchmark datasets: Set5 [6], Set14 [51], BSDS100 [28] (all three containing natural images of low to medium resolution), Urban100 [18] (high-resolution urban scenes), and Manga109 [29] (high-resolution Japanese manga illustrations). Like prior work, we use peak-signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric, where PSNR is computed in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets. Additional metrics are given in Appendix 0.C. Not all numbers could be computed for competing methods, because some do not provide public code or checkpoints, see Appendix 0.E.
Backbones. Following related work, we combine our method with three standard backbones for super-resolution and image restoration: (i) EDSR-baseline [24] (1.22 M parameters); (ii) RDN [52] (22.0 M parameters); and (iii) SwinIR [23] (12.0 M parameters).
@@
EDSR- baseline [24] (1.22 M) + interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
MetaSR [17] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
@@ 4.1 Super-Resolution Performance
We start by evaluating the three variants of our method (Thera-S, -M, and -L) on the hold-out DIV2K validation set, following the setup described above. Table 1 shows the PSNR values obtained by all investigated methods, for both in-distribution (×$\times 2$×$\times 4$) and out-of-distribution (×$\times 6$×$\times 30$) scaling factors. Thera achieves state-of-the-art performance. Thera-L outperforms all previous methods on 21 out of 24 cases, often by a considerable margin (e.g., $\times$3 using EDSR-baseline). Thera-M, which adds 3$\times$ fewer parameters to the backbone, still performs on par with or better than competitors that have comparable parameter counts. For scaling factors higher than ×$\times 2$, Thera-M also matches the performance of the more expensive, attention-based methods CiaoSR and CLIT, despite requiring 30×30$\times$30 × fewer parameters than CLIT.
Notably, even our minimal version Thera-S, with the same anti-aliasing guarantees but with only 8,256 parameters on top of the backbone, reaches very competitive performance. It mostly outperforms all but the most parameter-intensive methods with the EDSR backbone. Using the more powerful SwinIR backbone, Thera-S outperforms all previous methods, including attention-based ones, on 6 out of 8 scaling factors, despite adding almost 2000×2000$\times$2000 × fewer learnable parameters than CLIT (respectively 170×170$\times$170 × fewer than CiaoSR). Similarly, with RDN our smallest variant performs best in 7 of 8 scaling factors. We see this as rather strong evidence for the power of the principled multi-scale observation model that is baked into our network; especially for higher scaling factors.
For a comparison with conventional interpolation, we also report numbers using Lanczos (sinc) resampling with the respective ×$\times 4$ backbone. This baseline is almost always outperformed by dedicated ASSR methods, indicating that the latter do learn scale-specific priors.
@@
We further report the performance of Thera on five popular benchmark datasets in Tab. 2. For these experiments, we use the SwinIR backbone. Overall, due to the unreasonably small data volume and out-of-distribution nature of some legacy datasets (e.g., 5 small images for Set5), the results are less consistent. We posit that future work on ASSR would benefit from switching to a larger, statistically more meaningful test bed. Nonetheless, our methods perform best in 12 out of 15 cases. We once again observe that the performance of Thera-S is often comparable to that of methods with orders of magnitude higher parameter counts.
When visually inspecting the super-resolution results – for examples see Fig. 5 – we observe that Thera achieves a more faithful and perceptually more convincing reconstruction, particularly in the presence of repeating structures. This again highlights the advantage of hard-wiring a theoretically principled sampling model that rules out signal aliasing. For further qualitative results see Appendix 0.B.
@@
LR input LIIF [11] LTE [22] CiaoSR [8] Thera-L (ours) GT
Figure 5: Qualitative examples from four datasets for $\times$6 super-resolution. We use the SwinIR [23] backbone for all methods. Differences between methods are highlighted with bounding boxes. More examples are given in Appendix 0.B.
@@ 4.2 Multi-Scale Consistency
We are interested in learning a super-resolution model that is consistent across scales. For instance, if we super-resolve an image with a scaling factor $\times$4 then downsample it by a factor $\times$2, we expect the result to match what we get by directly super-resolving the original image by a factor $\times$2. We report in Tab. 3 the quality of this match for several pairs of scaling factors, both for Thera with its in-built observation model and for three comparison methods where the observation model must be discovered from data. First, we notice that higher super-resolution quality does not correlate with better scale consistency, e.g., LTE in most cases achieves better consistency than CiaoSR. Second, while other methods generally exhibit good consistency, even for out-of-distribution (OOD) scales, our formal model is superior for all scaling factor combinations, often by a large margin, while requiring only a single parameter $\kappa$ to be learned (if any, since $\kappa$ can also be fixed as we show below).
Table 3: Multi-scale consistency (in PSNR) of super-resolution methods across various pairs of scaling factors.
@@ 4.3 Ablation Studies
In Tab. 4, we ablate individual components and design choices of our method to understand their contributions to overall performance. These comparisons use Thera-M with EDSR-baseline backbone, but are also representative of other settings.
Single scale training. We ran three experiments using a single scale ($\times$2, $\times$3, or $\times$4) to test how this affects scale generalization. $\kappa$ was fixed for these experiments as multi-scale training is required to optimize it. As expected, we observe equal or even superior performance compared of single-scale training when tested at the training scale (marked in yellow in Tab. 4), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
Trainable $\bm{\kappa}$. Fixing $\kappa$ at the theoretically derived value (Eq. 5) leads to a small drop in performance. This suggests that there remain effects that are not accounted for by our proposed observation model, but that these are very minor.
Geometric self-ensemble. In line with previous work [42, 24, 43] we see a notable performance boost with geometric self-ensembling (GSE). Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
Total variation. The regularizer has a negligible effect for in-domain scaling factors, but performance degrades significantly without it for out-of-distribution scales.
No thermal activations. We replace thermal activations (Eq. 4) with standard ReLU activations. What remains is only the hypernetwork controlling the parameters of the local fields. A consistent loss in performance shows the impact of the proposed thermal activation underlying our multi-scale representation.
Shared components. We observe a noticeable performance drop, at all scales, if we make the components $\mathbf{W}_{1}$ an output of the hypernetwork $\Psi$. Despite the greater flexibility afforded by individual per-field components, it seems challenging to learn them jointly with other, correlated parameters.
Table 4: Ablation experiments with Thera-M using the EDSR-baseline backbone.
@@ 4.4 Limitations and Future Work
Thermal activations as introduced in this paper, and by extension our proposed ASSR method, come with relatively strict architectural requirements that only allow for a single hidden layer in the neural field. While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder. We have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, and we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work. This could result in even higher parameter efficiency, and potentially better generalization. It may also be possible to construct similar formulations that do not require the usage of a hypernetwork, which could enable more streamlined training.
Relative to other methods, Thera’s performance slightly drops for $\times$2 super-resolution when compared to higher scaling factors. We speculate that at small scaling factors Thera’s isotropic, continuous field representation is increasingly challenged by anisotropic effects from the oriented, discrete pixel grid. Further explorations are necessary to better understand this transition regime and perhaps come up with hybrid representations (which could potentially benefit also other methods at the interface of gridded and implicit representations like Instant-NGP [33]).
Finally, we expect that more advanced priors than TV could be even more effective at regularizing $\Phi$. Priors at $t=0$, made possible by Thera, have the potential to regularize the continuous signal itself, and therefore improve SR performance for all scaling factors.